What is 1/6 of 300?

Answer by Amrita Battula(2)


1/6 of 300 = ?
=> (300 * 1)/6 = 300/6 = 50 (since, 6 clearly divides 300 with zero remainder)
similarly, 1/5 of 250 is 250/5 = 50